OpenStudy (anonymous):

LIMITS-Seems easy but somehow I don't know what to do with the fractions Find the limit(if it exists) as x approaches 0 for the function (1/(X+1)-1)/x .

OpenStudy (psymon):

\[\frac{ 1 }{ \frac{ (x+1)-1 }{ x } }\] Something like that? Can't quite tell how you meant to type it up/

OpenStudy (anonymous):

Oops let me try again. ((1/(x+1))-1)/x

OpenStudy (psymon):

\[\frac{ 1 }{ x+1 }-\frac{ 1 }{ x }\] correct now?

OpenStudy (psymon):

Oh, I think I know what you meant to type up: \[\frac{ \frac{ 1 }{ x+1 } -1}{ x }\]

OpenStudy (anonymous):

no the first fraction is subtrated by 1, then that whole quantity is divided by x. oh yeah thats it c:

OpenStudy (psymon):

Okay, so just making sure we do our algebra right then. Alright, so this is how Id do it. Id make that top portion all into one fraction by making a common denominator. Basically, id take that -1 and multiply top and bottom by x+1, which allows me to make it into one fraction like this: \[\frac{ \frac{ 1-(x+1) }{ x+1 } }{ x }\implies \frac{ \frac{ -x }{ x+1 } }{ x }\] Now this is just the division of two fractions. The first fraction is -x/(x+1) and the second fraction is x/1. So I can rewrite what we have like this: \[\frac{ -x }{ x+1 }\div \frac{ x }{ 1 }\implies \frac{ -x }{ x+1 }*\frac{ 1 }{ x }=\frac{ -x }{ x(x+1) }\] So of course the x on tiop and bottom cancel out to leave \[\frac{ -1 }{ x+1 } \]Now apply your limit : )

OpenStudy (anonymous):

Aha! I can see it clearly now. The limit is -1.

OpenStudy (psymon):

yep yep :3 Honestly, its often not the calculus that sucks, its the algebra, lol.

OpenStudy (anonymous):

Haha I can see it coming back to haunt me already. Anyway that was superb, thanks for your time. :D

OpenStudy (psymon):

No problem : )